### v1

# Translators Asymptotic to Planes

Abstract.
We prove that a vertical plane is the only complete translator, properly immersed in $\mathbb{R}^{3}$ and having finite topology, whose ends are asymptotic to vertical planes.
1. Introduction
Translators arise naturally in the study of mean curvature flow as self-similar solutions that move by translation. Their geometric and analytic properties have been studied extensively, see for instance [1, Chapter 13], mainly because of their role in understanding type-II singularities of mean curvature flow [3, 4].
Our main theorem classifies complete properly immersed translators with finite topology whose ends are asymptotic to vertical planes, see Definition $2.$
Theorem $1.$
Let $\Sigma$ be a complete translator properly immersed in $\mathbb{R}^{3}$ with finite topology whose ends are asymptotic to vertical planes. Then $\Sigma$ is a vertical plane.
Our study was inspired by Khan’s work on complete embedded translators with finite total curvature [5], which, among other things, shows that such surfaces satisfy the hypotheses of Theorem $1.$ We therefore have the following corollary of Theorem $1.$
Corollary 1.2.
Every complete translator embedded in $\mathbb{R}^{3}$ with finite total curvature is a vertical plane.
Our proof of Theorem 1.1 relies principally on an identity expressing the total mean curvature of any smooth compact translator as the flux of $e_{3}^{\top}$ across its boundary, see Lemma 2.1.
2. Preliminaries
A translator $\Sigma\subset\mathbb{R}^{3}$ is a smooth surface whose mean curvature vector $H$ satisfies
$H=e_{3}^{\perp}.$
Equivalently, the family of surfaces $\Sigma+te_{3}$ for $t\in\mathbb{R}$ is a solution to the mean curvature flow.
In the following lemma, we express the total mean curvature of any smooth compact translator as the flux of $e_{3}^{\top}$ across its boundary.
Lemma 2.1.
For a smooth compact translator $\Sigma$ with boundary $\partial\Sigma$ we have
(1) $\int_{\Sigma}|H|^{2}=\int_{\partial\Sigma}\langle\eta,e_{3}^{\top}\rangle,$
where $\eta$ is the outward conormal to $\partial\Sigma$.
Proof.
Integrate the identity
$\operatorname{div}_{\Sigma}e_{3}^{\top}=-\operatorname{div}_{\Sigma}e_{3}^{\perp}=\langle H,e_{3}^{\perp}\rangle=|H|^{2}$
and apply the divergence theorem. ∎
Finally, we define what it means for an end to be asymptotic to a vertical plane.
Definition $2.$
Let $\Sigma_{i}$ be an end of a translator $\Sigma$, i.e. a component of $\Sigma\setminus B_{r}(0)$ for some sufficiently large radius $r$. We say that $\Sigma_{i}$ is asymptotic to a vertical plane if there exists a vertical plane $\Pi_{i}$ such that the following holds:
$\Sigma_{i}=\operatorname{graph}u_{i}$ for a function $u_{i}:\Pi_{i}\setminus\Omega_{i}\to\mathbb{R}$, where $\Omega_{i}\subset\Pi_{i}$ is open and bounded, and $\|u_{i}\|_{C^{1}(\Pi_{i}\setminus B_{R}(0))}\to 0$ as $R\to\infty$.
3. Proof of Theorem 1.1
To prove Theorem 1.1 we produce an exhaustion of $\Sigma$ by smooth, compact regions such that the boundary integral in (1) converges to zero. This implies that $H$ vanishes identically and hence $\Sigma$ is a vertical plane.
We first prove a variant of Theorem 1.1 which imposes a slightly different hypothesis on the ends.
Proposition 3.1.
Let $\Sigma$ be a complete translator properly immersed in $\mathbb{R}^{3}$. Suppose there exists a radius $r>0$ such that:
  • $\Sigma\setminus B_{r}(0)$ has finitely many components, denoted $\Sigma_{i}$;
  • each component $\Sigma_{i}=\operatorname{graph}u_{i}$ for a function $u_{i}:\Pi_{i}\setminus\Omega_{i}\to\mathbb{R}$, where $\Omega_{i}\subset\Pi_{i}$ is open and bounded, and
    (2) $\lim_{R\to\infty}\int_{\Pi_{i}\,\cap\,\partial B_{R}(p_{i})}|Du_{i}|^{2}=0$
    for some $p_{i}\in\Pi_{i}$.
Then $\Sigma$ is a vertical plane.
Proof.
For each $\Sigma_{i}$, up to an ambient isometry we may assume $\Pi_{i}=\{x_{1}=0\}$ and $p_{i}=0$. We introduce polar coordinates for $\Pi_{i}$, conflating $(0,r\cos\theta,r\sin\theta)$ with $(r,\theta)$. We may then express points in $\Sigma_{i}$ as $(u_{i}(r,\theta),r,\theta)$. For all sufficiently large $R$ we define map
$\gamma_{i,R}(\theta):=(u_{i}(R,\theta),R,\theta),$
giving rise to a smoothly embedded closed curve
$\Gamma_{i,R}:=\gamma_{i,R}(S^{1})\subset\Sigma_{i}.$
Since there are only finitely many components $\Sigma_{i}$ and $\Sigma$ is properly immersed, for each large $R$ the collection of curves $\Gamma_{i,R}$ bounds a compact region in $\Sigma$, which we denote by $\Sigma_{R}$. We will demonstrate that
$\lim_{R\to\infty}\int_{\Gamma_{R,i}}\langle\eta,e_{3}^{\top}\rangle=0$
for each $i$, where $\eta$ is the outward unit conormal to $\partial\Sigma_{R}$. The claim then follows, since Lemma 2.1 implies that
$\lim_{R\to\infty}\int_{\Sigma_{R}}|H|^{2}=0$
and hence $H$ vanishes everywhere, meaning $\Sigma$ can only be a vertical plane.
For the following computation let $i$ and $R$ be fixed and to ease notation let us write $\gamma$ instead of $\gamma_{i,R}$, $\Gamma$ instead of $\Gamma_{i,R}$, etc. The outward conormal to $\Gamma$ is
$\eta(R,\theta)=\frac{\gamma^{\prime}(\theta)}{|\gamma^{\prime}(\theta)|}\times\nu(R,\theta).$
where $\nu$ is the upward unit normal to $\Sigma$, given by
$\nu=\frac{e_{1}-Du}{\sqrt{1+|Du|^{2}}}.$
We therefore have
$\int_{\Gamma}\langle\eta,e_{3}^{\top}\rangle$ $=\int_{0}^{2\pi}\left\langle\frac{\gamma^{\prime}(\theta)}{|\gamma^{\prime}(\theta)|}\times\nu(R,\theta),e_{3}\right\rangle|\gamma^{\prime}(\theta)|\,d\theta$
$=\int_{0}^{2\pi}\langle\gamma^{\prime}(\theta)\times\nu(R,\theta),e_{3}\rangle\,d\theta.$
Evaluating the integrand at $(R,\theta)$ gives
$\langle\gamma^{\prime}(\theta)\times\nu,e_{3}\rangle$ $=\langle\gamma^{\prime}(\theta),e_{1}\rangle\langle\nu,e_{2}\rangle-\langle\gamma^{\prime}(\theta),e_{2}\rangle\langle\nu,e_{1}\rangle$
$=\frac{1}{\sqrt{1+|Du|^{2}}}(R\sin\theta|D_{2}u|^{2}-R\cos\theta D_{3}uD_{2}u+R\sin\theta)$
$=R\sin\theta+R\cdot O(|Du(R,\theta)|^{2}),$
and hence we have
$\int_{\Gamma}\langle\eta,e_{3}^{\top}\rangle$ $=\int_{0}^{2\pi}R\sin\theta\,d\theta+\int_{0}^{2\pi}R\cdot O(|Du(R,\theta)|^{2})\,d\theta$
$=\int_{0}^{2\pi}R\cdot O(|Du(R,\theta)|^{2})\,d\theta.$
The hypothesis (2), expressed in polar coordinates, ensures that the right hand side tends to zero as $R\to\infty$. This completes the proof. ∎
Our aim now is to prove that any end which is asymptotic to a vertical plane automatically satisfies (2). This will make use of the following interior gradient estimate, which is obtained by a standard application of Bernstein’s technique. [^1] Møller pointed out to us that the same estimate can be obtained by the different arguments contained in [2, Proposition 3.10] and a standard interpolation inequality.
Lemma 3.2.
Let $\Sigma$ be a translator of the form $\operatorname{graph}u$, where $u:\Pi\cap\overline{B_{R}(p)}\to\mathbb{R}$, $\Pi$ is a vertical plane containing the point $p$, and we assume $R\geq 1$. If $|Du|^{2}\leq\frac{1}{4}$ holds everywhere in $\Pi\cap\overline{B_{R}(p)}$ then we have
$|Du(p)|^{2}\leq\frac{400}{R}\max_{\Pi\,\cap\,\partial B_{R}(p)}u^{2}.$
Proof.
Up to an ambient isometry we may assume $\Pi=\{x^{1}=0\}$. A unit normal to $\Sigma$ is given by
$\nu:=\frac{e_{1}-Du}{\sqrt{1+|Du|^{2}}},$
and we have
$\langle H,\nu\rangle=\bigg(\delta^{ij}-\frac{D^{i}uD^{j}u}{1+|Du|^{2}}\bigg)\frac{D_{i}D_{j}u}{\sqrt{1+|Du|^{2}}}=:a^{ij}\frac{D_{i}D_{j}u}{\sqrt{1+|Du|^{2}}}$
and
$\langle e_{3},\nu\rangle=-\frac{D_{3}u}{\sqrt{1+|Du|^{2}}}.$
Therefore, the translator condition for $\Sigma$ is equivalent to
$a^{ij}D_{i}D_{j}u+D_{3}u=0.$
Here and throughout the proof indices range over $\{2,3\}$.
Let us write $L:=a^{ij}D_{i}D_{j}+D_{3}$. We then have
$L\frac{u^{2}}{2}=a^{ij}D_{i}uD_{j}u=\frac{|Du|^{2}}{1+|Du|^{2}}.$
and, using the Cauchy–Schwarz inequality,
$L\frac{|Du|^{2}}{2}$ $=-D^{k}uD_{k}a^{ij}D_{i}D_{j}u+a^{ij}D_{i}D^{k}uD_{j}D_{k}u\geq\bigg(1-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)|D^{2}u|^{2}.$
Let $\varphi\geq 0$ be a cutoff function and $\Lambda>0$ a constant, both to be chosen later. From the above computations we see that the auxiliary quantity $G:=\varphi^{2}\frac{|Du|^{2}}{2}+\Lambda\frac{u^{2}}{2}$ satisfies
$LG$ $=\frac{|Du|^{2}}{2}L\varphi^{2}+a^{ij}D_{i}\varphi^{2}D_{j}|Du|^{2}+\varphi^{2}L\frac{|Du|^{2}}{2}+\Lambda L\frac{u^{2}}{2}$
$\geq\frac{|Du|^{2}}{2}L\varphi^{2}+a^{ij}D_{i}\varphi^{2}D_{j}|Du|^{2}$
$\qquad+\bigg(1-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)\varphi^{2}|D^{2}u|^{2}+\frac{\Lambda|Du|^{2}}{1+|Du|^{2}}.$
Using $|a|\leq 2$ we estimate
$L\varphi^{2}\geq-4\varphi|D^{2}\varphi|-4|D\varphi|^{2}-2\varphi|D_{3}\varphi|$
and
$a^{ij}D_{i}\varphi^{2}D_{j}|Du|^{2}$ $\geq-8\varphi|D\varphi||Du||D^{2}u|\geq-\frac{1}{2}\varphi^{2}|D^{2}u|^{2}-32|D\varphi|^{2}|Du|^{2},$
leading to
$LG$ $\geq-|Du|^{2}(2\varphi|D^{2}\varphi|+34|D\varphi|^{2}+\varphi|D_{3}\varphi|)$
$\qquad+\bigg(\frac{1}{2}-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)\varphi^{2}|D^{2}u|^{2}+\frac{\Lambda|Du|^{2}}{1+|Du|^{2}}.$
Let $\varphi$ be chosen so that $\varphi(p)=1$, $\varphi\equiv 0$ on $\partial B_{R}(p)$, $0\leq\varphi\leq 1$ and
$R|D\varphi|\leq 2,\qquad R^{2}|D^{2}\varphi|\leq 8.$
In addition, let $\Lambda=400R^{-1}$. Inserting these choices above and using $R\geq 1$, we finally arrive at
$LG$ $\geq\bigg(\frac{400}{1+|Du|^{2}}-200\Bigg)R^{-1}|Du|^{2}+\bigg(\frac{1}{2}-5\frac{|Du|^{2}}{1+|Du|^{2}}\bigg)\varphi^{2}|D^{2}u|^{2}.$
The right-hand side is nonnegative in $B_{R}(p)$ since we are assuming $|Du|^{2}\leq\frac{1}{4}$, so by the maximum principle
$\frac{|Du(p)|^{2}}{2}\leq\max_{B_{R}(p)}G=\max_{\partial B_{R}(p)}G=\frac{400}{R}\max_{\partial B_{R}(p)}\frac{u^{2}}{2}.$
We now complete the proof of our main theorem.
Proof of Theorem $1.$
Let $\Sigma\subset\mathbb{R}^{3}$ be a complete translator satisfying the hypotheses of Theorem $1.$ We proceed by verifying that each one of the finitely many annular ends satisfies equation (2), so that the claim follows from Proposition 3.1.
Consider an annular end $\Sigma_{i}=\operatorname{graph}u_{i}$ and fix a point $p_{i}\in\Pi_{i}$ for reference. Since, by hypothesis, we have
$\sup_{\Pi_{i}\setminus B_{R}(p_{i})}|Du_{i}|\to 0\qquad\text{as}\qquad R\to\infty,$
we certainly have $|Du_{i}|\leq 1/4$ in the region $\Pi_{i}\setminus B_{R/2}(p_{i})$ whenever $R$ is sufficiently large. Therefore, given any $q\in\Pi_{i}\setminus B_{R}(p_{i})$, by Lemma 3.2 we have
$|Du_{i}(q)|^{2}\leq\frac{800}{R}\max_{\partial B_{R/2}(q)}u_{i}^{2}$
for sufficently large $R$. Since $q\in\Pi_{i}\setminus B_{R}(p_{i})$ was chosen arbitrarily we conclude that
$\sup_{\Pi_{i}\setminus B_{R}(p_{i})}|Du_{i}|^{2}\leq\frac{800}{R}\sup_{\Pi_{i}\setminus B_{R/2}(p_{i})}u_{i}^{2}$
for sufficiently large $R$. But we also assume
$\sup_{\Pi_{i}\setminus B_{R/2}(p_{i})}u_{i}^{2}\to 0\qquad\text{as}\qquad R\to\infty,$
so this shows that
$\sup_{\Pi_{i}\setminus B_{R}(p_{i})}|Du_{i}|^{2}\leq o(R^{-1})$
and hence
$\int_{\Pi_{i}\,\cap\,\partial B_{R}(p_{i})}|Du_{i}|^{2}=\int_{0}^{2\pi}R|Du_{i}|^{2}\,d\theta=o(1)$
as $R\to\infty$, which is (2). ∎